OpenStudy (anonymous):

solve each system of equations 6w-4v=11 3w=23-5v

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

can u answer this ?

OpenStudy (anonymous):

6w-4v=11 3w+5v=23 --------- 6w-4v=11 -2(3w+5v=23) ------------ 6w-4v=11 -6w-10v=-46 ------------ -14v=-35 v=2.5 6w-4(2.5)=11 .... the original equation 6w-10=11 6w=21 w=3.5 v = 2.5, w = 3.5 check... 6(3.5)-4(2.5)=11 3(3.5)+5(2.5)=23

OpenStudy (anonymous):

Good?

OpenStudy (anonymous):

Can you please put that I gave you a good answer? Thanks.